OpenStudy (anonymous):

Solve the inequality 2x < 8 - 4(2 - x)

OpenStudy (jdoe0001):

how would you solve say the EQUALITY of -> 2x = 8 - 4(2 - x) ?

OpenStudy (anonymous):

well i would distribute the -4

OpenStudy (jdoe0001):

right, so... what would you end up with?

OpenStudy (anonymous):

well i tried solving it myself and i got 4 but my teachersaid it was wrong so then i got 2

OpenStudy (anonymous):

First, let's work on the left hand side of your inequality, the 2x This means, for instance, to see if it can be simplified at all. Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x So, all-in-all, the left hand side of your inequality can be written as: 2x Now, let's work on the right hand side of your inequality, the 8 - 4(2 - x) 2-x evaluates to 2-x Multiply 4 by 2-x we multiply 4 by each term in 2-x term by term. This is the distributive property of multiplication. Multiply 4 and 2 1 4 × 2 = 8 Multiply 4 and -x Multiply 1 and x The x just gets copied along. x 4 × -x = -4x 4*(2-x) evaluates to 8-4x 8 - 8 = 0 The answer is 4x 8-4*(2-x) evaluates to 4x The right hand side of your inequality can be written as: 4x So with these (any) simplifications, the inequality we'll set out to solve is: 2x ‹ 4x Move the 4x to the left hand side by subtracting 4x from both sides, like this: From the left hand side: 2x - 4x = -2x The answer is -2x From the right hand side: 4x - 4x = 0 The answer is 0 Now, the inequality reads: -2x ‹ 0 To isolate the x, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the x on the left side of the inequality. The last step is to divide both sides of the inequality by -2 like this: To divide x by 1 The x just gets copied along in the numerator. The answer is x -2x ÷ -2 = x 0 ÷ -2 = 0 Note! Here's an important point about inequalities: We are now dividing by -2, which is a number less than zero! This changes the form of your inequality. Or, in other words, your ‹ will now become a ›. The solution to your inequality is: x › 0 So, your solution is: x must be greater than 0

OpenStudy (anonymous):

thank you so much